# Molecular clouds: do they deserve a non-Gaussian description?

### v1

@@
11institutetext: Laboratoire de Physique de l’École normale supérieure, ENS, Université PSL, CNRS, Sorbonne Université, Université Paris Cité, Paris, France 22institutetext: Observatoire de Paris, Université PSL, Sorbonne Université, LERMA, CNRS UMR 8112, 75014 Paris, France
Molecular clouds: do they deserve a non-Gaussian description?
@@
Pablo Richard Molecular clouds: do they deserve a non-Gaussian description?Molecular clouds: do they deserve a non-Gaussian description?    Erwan Allys Molecular clouds: do they deserve a non-Gaussian description?Molecular clouds: do they deserve a non-Gaussian description?    François Levrier Molecular clouds: do they deserve a non-Gaussian description?Molecular clouds: do they deserve a non-Gaussian description?    Antoine Gusdorf Molecular clouds: do they deserve a non-Gaussian description?Molecular clouds: do they deserve a non-Gaussian description?Molecular clouds: do they deserve a non-Gaussian description?Molecular clouds: do they deserve a non-Gaussian description?    Constant Auclair Molecular clouds: do they deserve a non-Gaussian description?Molecular clouds: do they deserve a non-Gaussian description?
(Submitted to A&A on July 13, 2024.)
@@
Molecular clouds are astrophysical objects whose complex non-linear dynamics is reflected in their complex morphological features. Many studies, investigating the bridge between higher-order statistics and physical properties, have shown the interest provided by non-Gaussian morphological features to grasp physical information. Yet, as this bridge is usually characterized in the supervised world of simulations, transferring it onto observations can be hazardous, especially when the discrepancy between simulations and observations remains unknown. In this paper, we aim at evaluating, directly from the observation data, the discriminating ability of a set of statistics. To do so, we develop a test that allows to compare, for a given unlabeled dataset, the informative power of two sets of summary statistics, relying on a notion of statistical compatibility. Contrary to supervised approaches, this test does not require the knowledge of any class or parameter corresponding to the data, but focuses instead on comparing the degeneracy levels of these descriptors, relying on a notion of statistical compatibility. We apply this test to column density maps of 14 nearby molecular clouds observed by Herschel, and iteratively compare different sets of usual summary statistics. We show that a standard Gaussian description of these clouds is highly degenerate but can be substantially improved when being estimated on the logarithm of the maps. This illustrates that low-order statistics, properly used, remain a very powerful tool. We then further show that such descriptions still exhibit a small quantity of degeneracies, some of which are lifted by the higher order statistics provided by reduced wavelet scattering transforms. This property of observations quantitatively differs from state-of-the-art simulations of dense molecular cloud collapse and is not reproduced by log fractional Brownian motion models. Finally we show how the summary statistics identified can be cooperatively used to build a morphological distance, which is evaluated visually, and gives very satisfactory results.
@@ 1 Introduction
Molecular clouds (MCs) play a key role in star formation. Their highly non-linear dynamics tends to couple spatial scales over a wide range, to create dense filamentary structures, in which clumps form that will eventually lead to prestellar cores (McKee & Ostriker, 2007). However, the precise role of each physical ingredient (such as turbulence, magnetic fields, and gravity, to name but a few) in this dynamics still remains to be fully understood.
@@
Figure 1: Column density map of a region in Perseus (top row) and sample of a log-Fractional Brownian Motion field (bottom row). Each image has 256 $\times$ 512 pixels but is tiled into 2 complementary patches of size 256 $\times$ 256 on which the statistics shown in Fig. 2 are computed. While these one-point and two-point statistics (some being non-Gaussian) are clearly compatible, these two images have manifestly different morphologies.
@@
Figure 2: Statistical properties of the maps from Fig. 1: PDFs (top), power spectra (center), and power spectra of the logarithms of the maps (bottom). The dashed vertical line in the power spectra plots is located at the $18.2\arcsec$ resolution of the column density map from the observational dataset. Power spectra are apodized as explained in Sec. 4. We see that the logFBM process is compatible with this observed portion of MC for all these statistics (some of theme being non-Gaussian), while having a manifestly different morphology, as revealed in Fig. 1. All these statistics are thus degenerate for such a comparison.
@@
A key step towards such an understanding certainly resides in our ability to decipher, in the morphology of these structured clouds, some signature of the physical processes at play. For instance, the power-law shape of the power spectrum in emission maps may trace properties of the turbulence (Miville-Deschênes et al. (2007); Federrath & Klessen (2013)), or the shape of the probability distribution function (PDF) of column density maps (Vázquez-Semadeni et al., 1997; Hennebelle & Chabrier, 2008; Kainulainen et al., 2009; Federrath & Klessen, 2013; Schneider et al., 2022; Appel et al., 2022) may trace the impact of gravitational collapse or stellar feedback as it transitions from a log-normal shape to the development of a heavy tail.
@@
However, the strong non-linear interplay between the physical processes at work leads to the emergence of non-Gaussian features in interstellar structures, such as filaments. This has motivated numerous studies to capture physical information beyond one-point and two-point statistics. Many approaches have been investigated including, for instance, diagnostics of the phase coherence of the Fourier modes (Levrier et al., 2006; Burkhart & Lazarian, 2016), bispectrum (Burkhart et al., 2009), structure functions (Heyer & Brunt, 2004), dendograms (Goodman et al., 2009), multiscale segmentation (Robitaille et al., 2019), scattering transforms (Allys et al., 2019; Regaldo-Saint Blancard et al., 2020; Saydjari et al., 2021), and neural networks (Peek & Burkhart, 2019; Zavagno et al., 2023).
@@
If higher-order statistics represent an appealing way of refining the geometrical description of complex ISM structures, they are nevertheless subject to the following caveats:
@@
    They should not screen out the potential of low-order statistics, but should come as complementary diagnostics (Burkhart & Lazarian, 2016). Low-order descriptors should not be underestimated: it is not because the processes under study are highly non-Gaussian that low-order statistics are only marginally informative. Yet, in numerous studies, such easy-access information is not considered, or even intentionally discarded [^1] The mean and variance of the data may be normalized, the power spectra flattened, and histograms (non-linearly) equalized., to emphasize the contribution of high-order statistics.
@@
    Refining the statistical description with higher-order terms complicates the connection between the statistical descriptions and the physical processes and associated parameters. Such a connection is thus usually learned in the supervised world of simulations, but transferring it to observations can be hazardous, especially when their distance to simulations remains unknown (Falgarone et al., 2004; Peek & Burkhart, 2019).
Figure 3: Footprints of the Herschel Gould Belt Survey (HGBS) fields used in this study, overlaid on the total thermal dust intensity at $353\,\mathrm{GHz}$ from the GNILC (Remazeilles et al., 2011) variable resolution data of Planck (Planck Collaboration, 2020).
@@
In this paper, we aim at identifying what sets of summary statistics are relevant to characterize the diversity of observed molecular clouds, without relying on simulations nor prior knowledge. In such an unsupervised framework, various observations cannot be grouped a priori in a similar class, so one often has to work in a very low data regime, which restricts dramatically the range of tools that can be used. This unsupervised and prior-knowledge free framework excludes features that require significant tuning to extract information, such as neural networks. This priority shift, from theoretical informativeness to actual information retrievable in this framework, might reshuffle the cards among the various sets of summary statistics to consider. This leads us to wonder to what extent non-Gaussian statistics can actually bring meaningful contributions in a fully observation-driven and unsupervised framework.
@@
To answer this question, we rely on a dataset of column density maps constructed from a survey of nearby molecular clouds. Our approach is to identify the amount of degeneracy that a set of summary statistics can have on this dataset, i.e., finding a pair of maps in this dataset that are compatible according to these statistics, but actually have a different morphology. A simplified illustration of such a degeneracy is given in Figs. 1 and 2, where we compare an observed column density map with a realization of a field whose logarithm is sampled from a Fractional Brownian Motion, that is a specific type of Gaussian process with power-law type constrained power spectrum. In this example, the PDFs of these maps, their power spectra, and the power spectra of their logarithms, all seem to be compatible, even though the maps clearly have different structures. This illustration underlines the limitation of those statistics, but only for this specific comparison, where we compare one observation to a specific known model. However, in an unsupervised framework, when working solely with unlabeled observations, we will exhibit these limitations by introducing additional statistics that could be used to lift these degeneracies.
@@
In this work, we use such a degeneracy diagnostic on the observed molecular clouds dataset to progressively build a robust set of summary statistics. We start with basic statistics, such as the power spectrum and one-point statistics. We evidence strong limitations that we investigate and substantially mitigate by taking the logarithm of the maps beforehand. We evidence further but moderate limitations of this improved set, leveraging higher-order statistics through the reduced wavelet scattering transform (RWST). This study, initially carried out only on the observational dataset, is then extended to logFBM data, as well as to a set of magnetohydrodynamics (MHD) simulations intended to reproduce observations of dense star-forming molecular clouds. We conclude by discussing the relevance and limitations of the final set of statistics we have constructed, and show how it can be used to assess the distance, in a statistical sense, between different maps of the interstellar medium.
The structure of the paper is as follows:
@@
  • In section 2, we present the set of MCs considered in this work and the corresponding column density maps, as well as the numerical simulations and the logFBM models.
@@
  • We present in section 3 the diagnostic of statistical compatibility that we will rely on, and our general methodology.
  • We present in section 4 the sets of statistics that will be confronted.
@@
  • We apply this methodology in section 5 to confront these sets of statistics on observations, and from these results we design an informative and low-dimensional set of summary statistics $\phi_{\text{final}}$.
@@
  • We finally define, in section 6, a distance based on $\phi_{\text{final}}$ that allows us to compare datasets, such as observations and simulations.
@@ 2 Data
In this section, we present the main dataset that we use for this study: an ensemble of $\sim$550 molecular hydrogen ($\mathrm{H}_{2}$) column density maps from a survey of nearby molecular clouds (MCs). Three other datasets will be used in this paper:
  • a set of $\sim$230 total gas column density maps built from magnetohydrodynamical (MHD) numerical simulations of dense molecular clouds, classified into three subsets with varying values of the magnetic field, from null (hydro), to medium (MHD), and high (MHD high B). These simulations are state-of-the art attempts to reproduce the physics at play in the early stages of star formation, including self-gravity and decaying turbulence, and are therefore well adapted to compare to our observations of the HGBS clouds.
  • A set of $\sim$500 synthetic maps sampled from logFBM models, whose parameters reproduce the diversity of the observational dataset. This type of purely synthetic fields has already been extensively used to model the interstellar medium (see, e.g., Elmegreen, 2002; Brunt & Heyer, 2002; Miville-Deschênes et al., 2007; Levrier et al., 2018).
@@
More details about these sets are given in appendix A. All these maps have a size of 512 $\times$ 512 pixels.
@@ 2.1 Observations: column density maps from the HGBS
We focus our study on a set of MCs targeted by the Gould Belt Survey (HGBS) (André et al., 2010) with the Herschel Space Observatory (Pilbratt et al., 2010), whose footprints on the sky are shown in Fig. 3. The HGBS combines two main criteria we require for this work:
  • it sampled numerous MCs with a diversity of physical and environmental conditions, from diffuse and quiescent regions with no sign of star-formation activity such as the Polaris Flare (Heithausen & Thaddeus, 1990; André et al., 2010; Miville-Deschênes et al., 2010) to very dense and active ones such as Orion B (Schneider et al., 2013) or the Aquila Rift cloud (Könyves et al., 2015). Examples of molecular column density maps from the HGBS are shown in Fig. 4. We however emphasize that this survey is limited to local clouds (distances $d\leq$ 500  pc) and does not cover the whole range of conditions expected in Galactic molecular clouds [^2] A further study could target more distant, more massive star forming regions with interferometric observations. We postpone such a study to a future paper..
  • It imaged a broad range of scales, from the full cloud size ($\sim 10\,\,$pc, corresponding to a few degrees for these nearby clouds) down to the $\sim 0.1\,\,$pc scales of filaments (Arzoumanian et al., 2011; André et al., 2014), which are spatially resolved for the nearest clouds [^3] The typical resolutions range from $10\,\arcsec$ at 70 $\mu$m to $36\,\arcsec$ at 500 $\mu$m., thus covering more than 2 orders of magnitude in spatial scales. This allows us to perform an in-depth morphological analysis, based on a local description of multi-scale interactions.
We consider the high resolution (18 ″) column density maps of 14 regions, produced with the procedure described in Appendix A of Palmeirim et al. (2013) and publicly available from the Herschel Gould Belt Survey Archive [^4] http://www.herschel.fr/cea/gouldbelt/en/Phocea/Vie_des_labos/Ast/

### v2

# Comparing the morphology of molecular clouds without supervision

@@
11institutetext: Laboratoire de Physique de l’École normale supérieure, ENS, Université PSL, CNRS, Sorbonne Université, Université Paris Cité, 75005 Paris, France 22institutetext: Observatoire de Paris, Université PSL, Sorbonne Université, LERMA, CNRS UMR 8112, 75014 Paris, France
Comparing the morphology of molecular clouds without supervision
@@
Pablo Richard Comparing the morphology of molecular clouds without supervisionComparing the morphology of molecular clouds without supervision    Erwan Allys Comparing the morphology of molecular clouds without supervisionComparing the morphology of molecular clouds without supervision    François Levrier Comparing the morphology of molecular clouds without supervisionComparing the morphology of molecular clouds without supervision    Antoine Gusdorf Comparing the morphology of molecular clouds without supervisionComparing the morphology of molecular clouds without supervisionComparing the morphology of molecular clouds without supervisionComparing the morphology of molecular clouds without supervision    Constant Auclair Comparing the morphology of molecular clouds without supervisionComparing the morphology of molecular clouds without supervision
(Submitted to A&A on July 13, 2024 / Accepted March 4, 2025)
@@
Molecular clouds are astrophysical objects whose complex nonlinear dynamics are reflected in their complex morphological features. Many studies investigating the bridge between higher-order statistics and physical properties have highlighted the value of non-Gaussian morphological features in capturing physical information. Yet, as this bridge is usually characterized in the supervised world of simulations, transferring it to observations can be hazardous, especially when the discrepancy between simulations and observations remains unknown. In this paper, we aim to evaluate, directly from the observation data, the discriminating ability of a set of statistics. To do so, we developed a test that allowed us to compare the informative power of two sets of summary statistics for a given unlabeled dataset. Contrary to supervised approaches, this test does not require knowledge of any class label or parameter associated with the data. Instead, it evaluates and compares the degeneracy levels of the summary statistics based on a notion of statistical compatibility. We applied this test to column density maps of 14 nearby molecular clouds observed by Herschel and iteratively compared different sets of typical summary statistics. We show that a standard Gaussian description of these clouds is highly degenerate but can be substantially improved when being estimated on the logarithm of the maps. This illustrates that low-order statistics, when properly used, remain a very powerful tool. We further show that such descriptions still exhibit a small quantity of degeneracies, some of which are lifted by the higher-order statistics provided by reduced wavelet scattering transforms. These degeneracies quantitatively differ between observations and state-of-the-art simulations of dense molecular cloud collapse , and they are not present for log-fractional Brownian motion models. Finally, we show how the summary statistics identified can be cooperatively used to build a morphological distance, which is evaluated visually and gives convincing results.
@@ 1 Introduction
Molecular clouds (MCs) play a key role in star formation. Their highly nonlinear dynamics tend to couple spatial scales over a wide range, creating dense filamentary structures in which clumps form, and these eventually lead to prestellar cores (McKee & Ostriker, 2007). However, the precise role of each physical ingredient (such as turbulence, magnetic fields, and gravity, to name but a few) in these dynamics still remains to be fully understood.
@@
Figure 1: Column density map of a region in Perseus (top row) and sample of a log-fractional Brownian motion (logFBM) field (bottom row). Each image has $256\times 512$ pixels but is tiled into two complementary patches of size $256\times 256$ on which the statistics shown in Fig. 2 are computed. While these one-point and two-point statistics (some being non-Gaussian) are clearly compatible, these two images have manifestly different morphologies.
@@
Figure 2: Statistical properties of the maps from Fig. 1: PDFs (top), power spectra (center), and power spectra of the logarithms of the maps (bottom). The dashed vertical line in the power spectra plots is located at the $18.2\arcsec$ resolution of the column density map from the observational dataset. Power spectra are apodized as explained in Sect. 4. We see that the logFBM process is compatible with this observed portion of MC for all these statistics (some of theme being non-Gaussian), while having a manifestly different morphology, as revealed in Fig. 1. All these statistics are thus degenerate for such a comparison.
@@
A key step toward such an understanding certainly resides in the ability to decipher, in the morphology of these structured clouds, some signature of the physical processes at play. For instance, the power-law shape of the power spectrum (PS) in emission maps may trace properties of the turbulence (Miville-Deschênes et al. (2007); Federrath & Klessen (2013)), or the shape of the probability distribution function (PDF) of column density maps (Vázquez-Semadeni et al., 1997; Hennebelle & Chabrier, 2008; Kainulainen et al., 2009; Federrath & Klessen, 2013; Schneider et al., 2022; Appel et al., 2022) may trace the impact of gravitational collapse or stellar feedback as it transitions from a log-normal shape to the development of a heavy tail.
@@
However, the strong nonlinear interplay between the physical processes at work leads to the emergence of non-Gaussian structures, such as filaments, in the interstellar medium (ISM). This has motivated numerous studies to capture physical information beyond one-point and two-point statistics. Many approaches have been investigated, including diagnostics of the phase coherence of Fourier modes (Levrier et al., 2006; Burkhart & Lazarian, 2016), bispectrum (Burkhart et al., 2009), structure functions (Heyer & Brunt, 2004), dendograms (Goodman et al., 2009), multiscale segmentation (Robitaille et al., 2019), scattering transforms (Allys et al., 2019; Regaldo-Saint Blancard et al., 2020; Allys et al., 2020; Saydjari et al., 2021), and neural networks (Peek & Burkhart, 2019; Zavagno et al., 2023).
@@
If higher-order statistics represent an appealing way to refine the morphological description of complex ISM structures, they are nevertheless subject to the following caveats:
@@
    They should not undermine the potential of low-order statistics but rather be regarded as complementary diagnostics (Burkhart & Lazarian, 2016). Low-order descriptors should not be underestimated; it is not because the processes under study are highly non-Gaussian that low-order statistics are only marginally informative. Yet, in numerous studies, such easy-access information is not considered (at times, it is even intentionally discarded [^1] The mean and variance of the data may be normalized, the power spectra flattened, and histograms (nonlinearly) equalized.) in order to emphasize the contribution of higher-order statistics.
@@
    Refining the statistical description with higher-order terms complicates the connection between the statistical descriptions and the physical processes and associated parameters. Thus, such a connection is usually learned in the supervised world of simulations, typically via a Fisher analysis (Allys et al., 2020) or a classification task (Saydjari et al., 2021). But transferring this connection to observations can be hazardous, especially when their similarity to simulations remains unknown (Falgarone et al., 2004; Peek & Burkhart, 2019) . Indeed, the informativeness of a set of summary statistics can be highly dataset dependent (Nunes & Balding, 2010; Blum et al., 2013).
@@
In this paper, we aim to identify which sets of summary statistics are relevant to characterize the diversity of observed molecular clouds without relying on simulations or prior knowledge. In such an unsupervised framework, various observations cannot be grouped a priori in a similar class, so one often has to work in a very low data regime, which dramatically restricts the range of tools that can be used. This unsupervised and prior-knowledge-free framework excludes features that require significant tuning to extract information, such as neural networks. This priority shift from theoretical informativeness to actual information retrievable in this framework might reshuffle the cards regarding the various sets of summary statistics to consider. This led us to wonder to what extent non-Gaussian statistics can actually bring meaningful contributions in a fully observation-driven and unsupervised framework.
@@
To answer this question, we rely on a dataset of column density maps constructed from a survey of nearby molecular clouds. Our approach is to identify the amount of degeneracy that a set of summary statistics can have on this dataset, that is, finding a pair of maps in this dataset that are compatible according to this set of statistics but that actually have a different morphology. A simplified illustration of such a degeneracy is given in Figs. 1 and 2, where we compare an observed column density map with a realization of a field, called logFBM, whose logarithm is sampled from a fractional Brownian motion (FBM), which is a specific type of Gaussian process with a power-law PS. In this example, the PDFs of these maps, their power spectra, and the power spectra of their logarithms all appear compatible, even though the maps clearly have different structures. For instance, the Perseus cloud shows filamentary features that are absent in the logFBM sample. This illustration underlines the limitation of these statistics, but only for this specific comparison, where we compare one observation to a specific known model. However, in an unsupervised framework, when working solely with unlabeled observations, we exhibit these limitations by introducing additional statistics that could be used to lift these degeneracies.
@@
In this work, we use such a degeneracy diagnostic on the observed molecular clouds dataset to progressively build a set of summary statistics that allows us to characterize their diversity, even in a low data regime. We start with basic statistics, such as the PS and one-point statistics. We evidence strong limitations that we investigate and substantially mitigate by applying a logarithmic transformation to the maps beforehand. We exhibit further but moderate limitations of this improved set, leveraging higher-order statistics through the reduced wavelet scattering transform (RWST). This study, initially carried out only on the observational dataset, is then extended to logFBM data as well as to a set of magnetohydrodynamics (MHD) simulations intended to reproduce observations of dense star-forming molecular clouds. We conclude by discussing the relevance and limitations of the final set of statistics we have constructed and show how it can be used to assess the distance, in a statistical sense, between different maps of the ISM.
The paper is structured as follows:
@@
  • In Sect. 2, we present the set of MCs considered in this work and the corresponding column density maps as well as the numerical simulations and the logFBM models.
@@
  • We present in Sect. 3 the diagnostic of statistical compatibility on which we rely and our general methodology.
  • We present in Sect. 4 the sets of statistics that are confronted.
@@
  • We apply this methodology in Sect. 5 to confront these sets of statistics on observations, and from these results, we design an informative and low-dimensional set of summary statistics $\phi_{\text{final}}$.
@@
  • We finally define, in Sect. 6, a morphological distance based on $\phi_{\text{final}}$ that allows us to compare datasets, such as observations and simulations.
@@ 2 Data
In this section, we present the main dataset that we use for this study: an ensemble of $\sim$550 molecular hydrogen ($\mathrm{H}_{2}$) column density maps from a survey of nearby MCs. Three other datasets are used in this paper:
  • A set of $\sim$230 $\mathrm{H_{2}}$ column density maps built from magnetohydrodynamical numerical simulations of dense molecular clouds, classified into three subsets with varying values of the magnetic field, from null (”Hydro high res”), to medium (”MHD”), and high (”MHD high B”). These simulations are used in this work as an example of state-of-the-art attempts to reproduce the physics at play in the early stages of star formation, including self-gravity and decaying turbulence . As it is often the case in studies of complex multi-physics ISM systems such as molecular clouds, these simulations focus on, and account only for, a partial variety of the physical processes at stake. For instance the set of simulations chosen here do not include any form of stellar feedback, that is nevertheless present in some regions of our observational dataset.
  • A set of $\sim$500 synthetic maps sampled from logFBM models, whose parameters reproduce the diversity of the observational dataset. This type of purely synthetic fields has already been extensively used to model the ISM (see, e.g., Elmegreen, 2002; Brunt & Heyer, 2002; Miville-Deschênes et al., 2007; Levrier et al., 2018).
@@
More details about these sets are provided in Appendix A. All these maps have a size of $512\times 512$ pixels.
@@ 2.1 Observations: Column density maps from the HGBS
We focus our study on a set of MCs targeted by the Gould Belt Survey (HGBS) (André et al., 2010) with the Herschel Space Observatory (Pilbratt et al., 2010) , whose footprints on the sky are shown in Fig. 14. The HGBS fulfills the two main criteria we require for this work:
  • It sampled numerous MCs with a diversity of physical and environmental conditions, from diffuse and quiescent regions with no sign of star formation activity such as the Polaris Flare (Heithausen & Thaddeus, 1990; André et al., 2010; Miville-Deschênes et al., 2010) to very dense and active ones such as Orion B (Schneider et al., 2013) or the Aquila Rift cloud (Könyves et al., 2015). Examples of molecular column density maps from the HGBS are shown in Fig. 3. However, we emphasize that this survey is limited to local clouds (distances $d\leq$ 500  pc) and does not encompass the full range of conditions expected in Galactic molecular clouds [^2] A further study could target more distant, more massive star forming regions with interferometric observations. We postpone such a study to a future paper..
  • It imaged a broad range of scales, from the full cloud size ($\sim 10\,\,$pc, corresponding to a few degrees for these nearby clouds) down to the $\sim 0.1\,\,$pc scale of filaments (Arzoumanian et al., 2011; André et al., 2014), which is spatially resolved for the nearest clouds [^3] The typical resolutions range from $10\arcsec$ at 70 $\mu$m to $36\arcsec$ at 500 $\mu$m., thus covering more than two orders of magnitude in spatial scales. This allows us to perform an in-depth morphological analysis, based on a local description of multiscale interactions.
[…]
Table 1: Summary of the properties of the different MCs studied in this paper, as well as their division into patches at different sampling.
Region Dist. Coord. # patches with resolution References
$d$ (pc) $l$ (°) $b$ (°) Tot. 6″ 9″ 12″ 18″
Aquila 436 28 4 57 36 16 4 1 1
Chamaeleon I 210 296 -16 28 20 6 2 0 2
Chamaeleon II III 160 302 -16 43 32 8 3 0 2
Corona Australis 160 0 -19 44 34 7 3 0 2
IC 5146 750 94 -5 4 4 0 0 0 2
Lupus I 182 339 17 17 12 4 1 0 3
Musca 200 301 -9 6 6 0 0 0 4
Ophiuchus 139 353 17 61 42 12 6 1 5
Orion B 450 205 -14 88 61 19 7 1 2, 6
Perseus 285 159 -20 44 32 10 2 0 2
Pipe Nebula 180 359 6 23 18 5 0 0 2
Polaris Flare 400 123 26 50 36 9 4 1 2, 6
Serpens 436 31 4 46 33 9 4 0 1
Taurus 145 170 -16 40 30 6 4 0 7
[…]
(1) Ortiz-León et al. (2018); (2) Zucker et al. (2020); (3) Galli et al. (2013); (4) Knude & Hog (1998); (5) Mamajek (2008); (6) Schlafly et al. (2014); (7) Yan et al. (2019).
We consider the high resolution (18″) $\mathrm{H_{2}}$ column density maps of 14 regions, produced with the procedure described in appendix A of Palmeirim et al. (2013) and publicly available from the Herschel Gould Belt Survey Archive [^5] http://www.herschel.fr/cea/gouldbelt/en/Phocea/Vie_des_labos/Ast/